How would you divide $10,000 if the 1st person gets twice as much as the 2nd person and the 2nd person gets 3 timese as much as the third person?
How much does the 1st person get?
How much does the 2nd person get?
How much does the 3rd person get?
Thank You! :)

Answer by gonzo(654)

let x = amount first person gets.
then 2*x = amount second person gets.
then 3*x = amount third person gets.
-----------------------------------------
total amount that everybody gets is 1*x + 2*x + 3*x = 6*x
so 6*x = 10,000
dividing both sides of the equation by 6 gets
x = 10,000/6
-----------------------------------------
take the 10,000 and divide it by 6
answer is 10,000 / 6 = 1666 and 2/3.
x = 1666 and 2/3.
------------------------------------------
person 1 gets 1666 and 2/3
person 2 gets 3332 and 4/3
person 3 gets 4998 and 6/3
-----------------------------------------
10,000 should equal to the sum of these.
1666 + 3332 + 4998 = 9996
2/3 + 4/3 + 6/3 = (2+4+6)/3 = 12/3 = 4
9996 + 4 = 10,000
-----------------------------------------